word2:

Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4> 20.3° 10.2° 0.2° 30.3°

word2:

@ThisGirlPretty

ThisGirlPretty:

Welcome to questioncove! and I'm sorry, I'm not good in math v.v @Nnesha HELPPPP @Vocaloid @Hero

word2:

@Nnesha

word2:

Is it 10.2?

Nnesha:

here is the formula u use to find the angle \[\huge\rm \cos \theta = \frac{u \cdot v}{\left| u \right|\left| v \right| } \]

word2:

so i just plug in 6,1 and 7,-4 in u and v and thats the answer?

Nnesha:

i didn't get 10.2. can you post your work please?

Nnesha:

well first you have to find the dot product of u/v and then find the magnitude of each vector are you familiar with dot product and magnitude ?

word2:

No , i just started this lesson and im very confused

word2:

i think i know what the dot product is

Nnesha:

ok dot product is a method to multiply the vectors if the given vectors are \[\large\rm u= < a_1 , b_2> ~~~,~~v= <a_2,b_2>\]

word2:

i think its 46 , because 6*7 + -1* -4

Nnesha:

ok dot product is a method to multiply the vectors if the given vectors are \[\large\rm u= < a_1 , b_2> ~~~,~~v= <a_2,b_2>\] \

Nnesha:

yes that's right

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nnesha ok dot product is a method to multiply the vectors if the given vectors are \[\large\rm u= < a_1 , b_2> ~~~,~~v= <a_2,b_2>\] \(\color{#0cbb34}{\text{End of Quote}}\) this is what i was trying to post https://prnt.sc/g418y6

Nnesha:

and the magnitude will be \[\left| u \right| = \sqrt{a^2+b^2}\]

word2:

i'm confused , what numbers do i use to find the magnitude

Nnesha:

there are different ways to write a vector \[\large\rm u= \color{Red}{a}i +\color{blue}{b}j ~~~~or~~~~ u= <\color{red}{a},\color{blue}{ b}>\] \

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nnesha there are different ways to write a vector \[\large\rm u= \color{Red}{a}i +\color{blue}{b}j ~~~~or~~~~ u= <\color{red}{a},\color{blue}{ b}>\] \ \(\color{#0cbb34}{\text{End of Quote}}\) https://prnt.sc/g41cen

Nnesha:

so the magnitude of u vector is \[\ \left| u \right|= \sqrt{a^2+b^2} = \sqrt{(6)^2+(-1)^2}\]

word2:

\[\sqrt{37}\]

word2:

so the magnitude is square root 37?

word2:

do i do the same for v?

Nnesha:

yes right and yes

word2:

\[\sqrt{(7)^2+(3)^2 } = \sqrt{58}\]

Nnesha:

it's -4 not 3 so that should be sqrt{65}

word2:

oh your right

word2:

what do i do with \[\sqrt{37} \sqrt{65}\]

Nnesha:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Nnesha here is the formula u use to find the angle \[\huge\rm \cos \theta = \frac{u \cdot v}{\left| u \right|\left| v \right| } \] \(\color{#0cbb34}{\text{End of Quote}}\) now u can plugin the values the numerator (dot product ) = 46 denominator (magnitude of u and v) = sqrt{65} times sqrt{37}

word2:

0.93799440

word2:

that's what i got when i plugged it into my calculator

Nnesha:

\[\cos \theta = 0.93799440 \] take inverse cos to find theta

word2:

\[\theta= 0.591406\]

Nnesha:

\[\large\rm \theta = \cos^{-1} (0.93799440)\]

Nnesha:

use the calculator^

word2:

\[\theta= 0.353998\]

Nnesha:

looks right

word2:

But thats not one of the answers

Nnesha:

do you have the answer choices ?? maybe they want the answer in degree not radians ?

word2:

20.3° 10.2° 0.2° 30.3°

Nnesha:

ohh okay so they want the answer in degree not radians in order to convert the answer to degree multiply it by 180/pi

word2:

20.282591356

word2:

when rounded its A , correct?

Nnesha:

yes

word2:

THANK YOU SO MUCH

Nnesha:

yw :=)) o^_^o

ThisGirlPretty:

Good job Nnesha and word2 :)!!